A simplified economy is specified as follows:

A. Goods market, all values C, I, G and NX values are in billions of C$:

Consumption Expenditure: C = 130 + 0.7(Y-T)Investment Expenditure: I = 1,300 - 530iGovernment Expenditure: G = 320Lump-sum Constant Taxes: T = 320Exports: 70Imports: 10


B. Money market, all Md values are in billions of C$:

Interest Rate: i = 0.07 or 7%Money Demand: Md = 710 - 2,000i


Note: Please keep your answers accurate to two decimal places.

a) Given the above information, solve for the following: the equilibrium Y, the money supply M, the consumption expenditure C, and the investment expenditure I.

Find Y, M,C, and I

Please help
1
Expert's answer
2016-11-09T09:30:08-0500
A. Goods market, all values C, I, G and NX values are in billions of C$: C = 130 + 0.7(Y - T), I = 1,300 - 530i, G = 320, T = 320, X = 70, M = 10.
B. Money market, all Md values are in billions of C$: i = 0.07 or 7%, Md = 710 - 2,000i.
a) The equilibrium Y is:
Y = C + I + G + NX = 130 + 0.7(Y - 320) + 1300 - 530*0.07 + 320 + 70 - 10,
0.3Y = 1,996.9,
Y = $6,656.33 billion.
The money supply Ms in equilibrium equals money demand Md, so:
Ms = Md = 710 - 2,000*0.07 = $570 billion.
The consumption expenditure C equals:
C = 130 + 0.7(Y - T) = 130 + 0.7*(6656.33 - 320) = $4,565.43 billion.
The investment expenditure I = 1300 - 530*0.07 = $1,262.9 billion